# Multiplexing in Networks and Diffusion

### v1

@@ 3.3. Multiplexing and Complex Diffusion
To keep the analysis as uncluttered as possible, we again focus on the case of two layers. We also consider a case where the correlation in transmission across layers is neither too high nor too low, so that there is an $\varepsilon$, to be determined in the proofs of the propositions below, for which $(1+\varepsilon)q_{A}q_{B}\geq f(2,\{A,B\})\geq q_{A}q_{B}$. Of course, a sufficient condition for this to hold is independent transmission. This condition is needed as with excessive positive or negative correlation in transmission, strange discrete behavior in transmission as a function of multiplexing can occur. [^19] For instance, if transmission is perfectly positively correlated, then one is always more likely to get two transmissions from a single multiplexed connection than two unmultiplexed connections, but is always more likely to get one transmission from the reverse. This then implies that the optimal configuration of connections depends on whether $\tau$ is even or odd, in complicated ways as a function of a node’s overall degrees in each layer. The restriction to two layers allows the results to highlight the more fundamental forces of multiplexing.
@@
There exist $10<\underline{\rho}<\overline{\rho}<$ 1 such that
@@
Consider a complex contagion ($\tau>1$), with nonnegative correlation in transmission across layers, so that $f(2,\{A,B\})\geq q_{A}q_{B}$, and two distributions such that $P^{\prime}\overline{\prec}P$ and both have positive steady-state infection rates. Also suppose that $\hat{D}_{A}+\hat{D}_{B}+2\hat{D}_{AB}>\tau$ for each $\hat{D}$ in the distribution $P$, so that each node has more than enough connections to become infected. There exist $10<\underline{\rho}<\overline{\rho}<$ 1 such that
@@ 3.4. Simulations
The theoretical results are based on the asymptotic statistical behavior of large random networks. To see how the results work in smaller empirical networks, we run simulations on the networks from the RCT villages. We simulate a Susceptible-Infected-Susceptible (SIS) diffusion process for the cases of both simple and complex diffusion and compare outcomes as multiplexing is varied. In order to compare across similar-sized networks where only multiplexing is changing, we take a given village network and construct two-layer networks by combining different pairs of empirical networks (which end up empirically having different multiplexing rates), in a way we specify below. We then perform many diffusion simulations on these two-layer networks for each village.
@@ 4. Concluding Discussion
Appendix A Proofs
@@
Proof of Proposition 2: Following the argument from the proof of Proposition 1, for any $\rho$ equation 3.4 has a higher solution for the less multiplexed type. Thus, starting with the steady state $\rho$ for the more multiplexed distribution, the new rates for all individuals are weakly and sometimes strictly higher for the less multiplexed distribution. This leads to a higher $\rho^{\prime}$. Iterating, this converges upward for all types to a limit which is the steady state. Conversely, if condition 3.1 is reversed, the convergence is downward for all types.
Proof of Proposition 3: It is enough to consider an individual $i$ with one change in their links where a multiplexed link to $j$ is then split to two neighbors $j,k$, each of which is connected to $i$ on a different layer and where $i$ was initially not connected to $k$. Our focus is on the pivotal cases:
@@
Supplementary Appendix:
Multiplexing in Networks and Diffusion
by Chandrasekhar, Chaudhary, Golub, Jackson
A.1. Supplementary Figures
Figure S1 plots the results from Table 2. $\hat{\beta}^{\ell}$ and both the 90% and 95% confidence intervals for each of the distinct layers are plotted. Seed centrality in the jati network is not statistically significantly associated with diffusion ($p=0.459$). Seed centrality in the advice, social, and kerorice networks all are significantly positively associated with diffusion. The point estimates are large, roughly a 59% increase.
@@ A.1. Supplementary Figures
Here we redo the analysis from Figure 8, but instead using the individual network data from Microfinance villages collected as part of Wave I of data collection in Banerjee et al. (2013) (instead of Wave II). [^20] We have individual-level gender-distinguished data in the Wave I network survey, which elicited links from $46$% of the households, giving us information on $70.84$% of the links.
A.2. Supplementary Tables
@@
A.3. Algorithms
@@
             Calculate $E_{i,t}=\sum_{j\in N}\sum_{l=1}^{2}G_{ij}^{(l)}\cdot\mathbb{I}(\sigma_{j,t}=I)\cdot q$;
@@
Step 1: Use keroricego, visitgo, and advice as the three matrices respectively;

### v2

@@ 4.3. Multiplexing and Complex Diffusion
To keep the analysis as uncluttered as possible, we again focus on the case of two layers. We also consider a case where the correlation in transmission across layers is nonnegative and not too high; formally,there is an $\varepsilon$, to be determined in the proofs of the propositions below, for which $(1+\varepsilon)q_{A}q_{B}\geq f(2,\{A,B\})\geq q_{A}q_{B}$. Of course, a sufficient condition for this to hold is independent transmission, under which $f(2,\{A,B\})=q_{A}q_{B}$. This condition is needed as with excessive positive or negative correlation in transmission, strange discrete behavior in transmission as a function of multiplexing can occur. [^14] For instance, if transmission is perfectly positively correlated, then one is always more likely to get two transmissions from a single multiplexed connection than two unmultiplexed connections, but is always more likely to get one transmission from the reverse. This then implies that the optimal configuration of connections depends on whether $\tau$ is even or odd, in complicated ways as a function of a node’s overall degrees in each layer. The restriction to two layers allows the results to highlight the more fundamental forces of multiplexing.
@@
There exist $0<\underline{\rho}<\overline{\rho}<1$ such that
@@
Consider a complex contagion ($\tau>1$), with nonnegative correlation in transmission across layers, so that $f(2,\{A,B\})\geq q_{A}q_{B}$, and two distributions such that $P^{\prime}\overline{\prec}P$ and both have positive steady-state infection rates. Also suppose that $\hat{D}_{A}+\hat{D}_{B}+2\hat{D}_{AB}>\tau$ for each $\hat{D}$ in the distribution $P$, so that each node has more than enough connections to become infected. There exist $0<\underline{\rho}<\overline{\rho}<1$ such that
@@
In terms of the reflection of the theory on the empirical results, the fact that more multiplexed villages had lower diffusion conditional on enough seeds to get the process going suggests that either the process was simple diffusion (which would be consistent with simply needing to know about the free phone giveaway), or that it had a relatively low threshold in most people and there was enough participation so that spreading links out was more likely to lead to a pivotal contact. As we do not fully observe the information spread in the villages, to get a deeper understanding of how the diffusion works in those villages we next do some simulations on those village networks to illustrate the theoretical results and explore some of the comparative statics.
@@ 4.4. Simulations
To get a deeper feeling for how the theory depends on the details of the process, we run simulations on the networks from the RCT villages. We simulate a Susceptible-Infected-Susceptible (SIS) diffusion process for the cases of both simple and complex diffusion and compare outcomes as multiplexing is varied. In order to compare across similar-sized networks where only multiplexing is changing, we take a given village network and construct two-layer networks by combining different pairs of empirical networks (which end up empirically having different multiplexing rates), in a way we specify below. We then perform many diffusion simulations on these two-layer networks for each village.
@@ 5. Discussion and Further Observations
Supplemental Online Appendix for
Multiplexing in Networks and Diffusion
by Chandrasekhar, Chaudhary, Golub, Jackson
Appendix A Additional Data Background
A.1. Descriptive Statistics
Descriptive statistics are presented in Table S1. The different layers exhibit significantly different patterns of connection. For example, in both datasets, the social layer is denser than the other layers and has among the highest levels of clustering. We observe a higher variance of node degrees in the decision layer than in other comparable layers (e.g., advice). [^15] In the RCT villages the social layer is significantly denser than kerorice, advice, or decision layers (p-values $0.009$, $0.000$, and $0.000$ respectively). The advice layer has significantly less variance (p-value $0.0006$) and more clustering (p-value $0.03$) than the decision layer. Similar patterns hold in the Microfinance villages.
Table S1. Descriptive Statistics
Network degree degree S.D. density triangles clustering
Microfinance villages
   social 15.296 7.841 0.079 2635.040 0.252
   kerorice 7.029 3.834 0.037 594.160 0.259
   advice 6.158 3.835 0.032 299.120 0.168
   decision 6.553 4.309 0.034 356.040 0.169
   money 8.512 5.036 0.044 681.960 0.193
   temple 1.709 1.899 0.009 52.040 0.175
   medic 6.530 3.911 0.034 369.400 0.188
   union link 75.428 32.542 0.368 314121.027 0.862
   intersect link 0.576 0.883 0.003 7.000 0.203
   jati 68.291 34.293 0.332 310150.907 1.000
RCT villages
   social 5.711 3.626 0.031 251.271 0.185
   kerorice 4.910 3.235 0.027 176.557 0.174
   advice 4.197 3.091 0.023 124.100 0.161
   decision 4.206 3.675 0.023 125.571 0.145
   union link 55.756 27.861 0.296 150771.400 0.913
   intersect link 1.812 1.829 0.010 38.871 0.229
   jati 52.633 28.599 0.279 150117.500 1.000
We also observe that the microfinance villages and RCT villages differ from each other in some descriptive statistics. Microfinance villages across all network layers are denser on average and exhibit higher levels of clustering, as can be seen in Table S1. The two samples also slightly differ in terms of village size. RCT villages have 197 households on average, while microfinance villages are larger with 216 households on average.
A.2. Additional RCT Details
In particular, they had to dial the provided promotional number and leave a “missed call.” This was a call that we registered but did not answer and was free for the participant to make, which was a standard technique for registration at the time. Registered callers were visited a few weeks later and received a reward. The individual rolled a pair of dice and received INR 25 $\times$ the number rolled. This yielded cash prizes of amounts ranging from INR 50 (for a 2) to INR 275 (for an 11). A roll of 12 was rewarded with a cell phone worth INR 3000. The expected value of the prize was INR 255, which was more than half of a day’s wage in the area.
A.2.1. Network Layers
In terms of notation, we define a multi-layered, undirected network for each village $v$, for layer $\ell=1,\ldots,L$, with $g_{ij,v}^{\ell}=1$ if either household $i$ or $j$ reported having a relationship of type $\ell$. We add another layer where $i$ and $j$ are linked if they belong to the same jati. For the Microfinance Village Sample, where GPS data are available, we construct a weighted graph where the $ij$ entry is the geographic distance between the two households.
We construct three synthetic network layers. The union layer has a link is present if a link exists in any layer. The intersection layer has a link is present if it exists in all layers. Finally, the total network is constructed as a weighted and directed network whose edge weights are the sums of indicators for links in all directed layers (using the raw directed nomination matrices, thus excluding jati and geography).
A.3. Principal Component Analysis (PCA)
We perform a principal component analysis with all of the layers (excluding the synthetic union, intersection layers and total network layers). We treat each pair of households (in a given village) as an observation, yielding $\sum_{v}\binom{n_{v}}{2}$ observations, where $n_{v}$ is the number of households in village $v$, and the number of dimensions is the number of layers $L$ in the given sample.
A.3.1. The Backbone Network Construction
The backbone network is built using the first $K$ principal components derived from the PCA. To select the optimal $K$ number of principal components the literature usually relied on a cutoff based on patterns of either decreasing eigenvalues or increasing variability of eigenvectors. Luo and Li (2016) combine these two approaches to better estimate the optimal $K$. They propose a new estimator, called the “ladle estimator” which minimizes an objective function that incorporates both the magnitude of eigenvalues and the bootstrap variability of eigenvectors. This approach exploits the pattern that when eigenvalues are close together, their corresponding eigenvectors tend to vary greatly, and when eigenvalues are far apart, the eigenvector variability tends to be small. By leveraging both sources of information, the ladle estimator can more precisely determine the rank of the matrix, and thus the optimal number of components to retain.
$w_{k}:=\lambda_{k}/\sum_{j=1}^{K}\lambda_{j}.$
For each village $v$, we then define a “backbone” network, $g^{\text{backbone}}$, from the principal components as a weighted graph where
$g_{ij,v}^{\text{backbone}}=Z_{ij,v}.$
A.4. Variable definitions
A.4.1. Diffusion Centrality
$DC_{j,v}^{\ell}:=\left[\sum_{t}^{T}(qg^{\ell}_{v})^{t}\cdot 1\right]_{j},$
$DC_{v}^{\ell}:=\sum_{j\in S_{v}}DC_{j,v}^{\ell}.$
A.4.2. The Multiplexing Score
We define a multiplexing score for household $i$ in village $v$ as
This measures the average fraction of relationship types household $i$ has with each of its neighbors. The numerator calculates the average number of links household $i$ has to each neighbor across all $L$ relationship types. It does this by first summing the number of links between household $i$ and each neighbor $j$ across all layers, dividing by the total number of layers $L$, and then summing this average across all neighbors $j$. The denominator counts the number of unique neighbors of household $i$ by summing an indicator for whether there is at least one link between $i$ and $j$ across any layer.
$\text{High Mpx}_{v}:={\bf 1}\left\{m_{v}>\text{median}(m_{1:v})\right\}.$
A.5. LASSO
To select a sparse set of network layers that best predict diffusion we perform a LASSO ($\ell_{1}$ penalized) regression. The regression of interest is given by
The regression of interest is given by
$y_{v}=\alpha+\sum_{\ell}\beta^{\ell}\cdot DC_{v}^{\ell}+X_{v}\Gamma+\epsilon_{v}$
where $DC_{v}^{\ell}$ represents the seed set diffusion centrality for layer $\ell$ in village $v$. We are interested in which $\beta^{\ell}$ are estimated to be non-zero and the consistent estimates of these parameters.
Given the high correlation between network layers, we fail to satisfy the irrepresentability condition which requires that the regressors of interest not be excessively correlated (Zhao and Yu, 2006). To overcome this problem, we use the Puffer transformation developed by Rohe (2015) and Jia et al. (2015), which recovers irrepresentability when the number of observations exceeds the number of variables. Although the regressors $(DC^{\ell}_{v})_{v,\ell}$ have correlated columns, by appropriately pre-conditioning the data matrix, we can force its columns to be orthogonal and therefore irrepresentable. Puffer-LASSO then recovers the set of relevant variables with probability tending to one exponentially fast in the number of observations, with consistent parameter estimates, that are asymptotically normally distributed with probability approaching one (Javanmard and Montanari, 2013; Jia et al., 2015; Taylor and Tibshirani, 2015; Lee et al., 2016; Banerjee et al., 2024a).
Appendix B Proofs
@@
Proof of Proposition 2: Following the argument from the proof of Proposition 1, for any $\rho$ equation 4.4 has a higher solution for the less multiplexed type. Thus, starting with the steady state $\rho$ for the more multiplexed distribution, the new rates for all individuals are weakly and sometimes strictly higher for the less multiplexed distribution. This leads to a higher $\rho^{\prime}$. Iterating, this converges upward for all types to a limit which is the steady state. Conversely, if condition 4.1 is reversed, the convergence is downward for all types.
Proof of Proposition 3: It is enough to consider an individual $i$ with one change in their links where one layer of a multiplexed link to $j$ is reassigned to a neighbor $k$, so that $i$ is connected to $j$ on one layer and to $k$ on a different layer, where $i$ was initially not connected to $k$. Our focus is on the pivotal cases:
@@
Appendix C Supplementary Figures
@@
Figure S1 plots the results from Table 1. $\hat{\beta}^{\ell}$ with both the 90% and 95% confidence intervals for each of the distinct layers. Seed centrality in the jati network is not statistically significantly associated with diffusion ($p=0.459$). Seed centrality in the advice, social, and kerorice networks all are significantly positively associated with diffusion. The point estimates are large, roughly a 59% increase.
@@
Here we redo the analysis from Figure 8, but instead we use the individual network data from Microfinance villages collected as part of Wave I of data collection in Banerjee et al. (2013) (instead of Wave II). [^17] We have individual-level gender-distinguished data in the Wave I network survey, which elicited links from 4646% of the households, giving us information on 70.8470.84% of the links.
C.1. Supplementary Tables
@@
  • Note: The “F-stat” and “p-val” columns correspond to the cumulative test comparing each specification with the intercept only benchmark. “F-stat marginal” and “p-val marginal” columns correspond to the marginal test when adding a given layer. Relative to Table 3 here we exclude the “constructed” intersection, union, total links, and backbone layers.
C.2. Algorithms
@@
      Calculate $E_{i,t}=\sum_{j\in N}\sum_{l=1}^{2}G_{ij}^{(l)}\cdot\mathbb{I}(\sigma_{j,t}=I)\cdot B_{ij,t}^{(l)}$, where $B_{ij,t}^{(l)}\sim\text{Bernoulli}(q)$ i.i.d.;
@@
1exStep 1: Use directed Kerorice, social, and advice as the three matrices respectively;